OpenStudy (anonymous):

Which is the simplified form of 18 times q squared times r to the fourth power all over 3 times q cubed times r.?

OpenStudy (anonymous):

@Alf3mb @bluescluesakaspongebo @johnweldon1993 @alex96 @rizwan_uet

OpenStudy (anonymous):

Is 6r^3/q the answer?

OpenStudy (anonymous):

Well I really don't know how to do it. I just assumed that 3/18 is 6, 4-3 =1, and 4-1=3.

OpenStudy (anonymous):

 *3-2=1